# Conveying Emotions to Robots through Touch and Sound

### v1

@@ 3.5 Consistency analysis
In order to transform feature values to a similar scale, ensuring all features contribute equally to the model, we standardized the tactile and audio features. After standardization, the numeric features of the dataset will have a mean of 0 and a standard deviation of 1. We then performed Principal Component Analysis (PCA) to reduce dimensionality and identify the key components capturing the most variance. The resulting principal component scores were integrated with the original emotion and participant identifiers. To evaluate the consistency of emotion conveyance, we calculated Intraclass Correlation Coefficients (ICC) for each emotion using the PCA scores and a two-way consistency model. As shown in Fig.3 and Table.1, all emotions demonstrated statistically significant consistency across participants, although some emotions exhibited comparatively low ICC values. The highest ICC was observed for “Attention”, indicating the most consistent expression, while “Surprise” had the lowest ICC, suggesting the greatest variability in expression.
@@
3.6 Variability analysis
The features do not follow a normal distribution. Therefore, we used permutational multivariate analysis of variance (PERMANOVA), which reveals a significant effect of emotion on the scaled combined features (). Pairwise comparisons with Holm correction show that most emotion pairs differ significantly and have consistent significance after p-value adjustment. However, there is no difference between calming and sadness (), comfort and sadness () based on touch and audio features, and there is no difference between confusion and comfort (), confusion and sadness (), disgust and sadness (), disgust and surprise (), happiness and surprise ().
Intuitively, it is expected that “surprise” is very similar to “happiness”. One explanation could be that both of the emotions belong to the high arousal zone. Similarly, there is no significant difference between “confusion” and “sadness”, or between “disgust” and “sadness”, as they belong to the low-arousal zone.
3.7 Decoding emotions
3.7.1 Data split
@@
3.7.2 Model comparison
We used various machine-learning techniques, including Random Forest, Support Vector Machine (SVM), k-Nearest Neighbors (k-NN), Gradient Boosting Machine (GBM), Decision Tree, and Naive Bayes (NB), to decode the emotions conveyed by touch and sound due to their robustness and proficiency in handling high-dimensional data. Six models were evaluated using 10-fold cross-validation to ensure robustness and accuracy. The results of these comparisons are presented in Table 2. In addition, we explored the most significant features; we leveraged feature importance scores from Random Forest and employed Recursive Feature Elimination (RFE) with SVM, identifying the top contributing features. The top three are Max Pressure, Mean Touch Duration and RMSE, which indicate that the features from both touch and audio channel are important.
| Models | RF | SVM | K-NN | GBM | DT | NB |
| Test accuracy | 38.33% | 40.00% | 28.89% | 35.00% | 31.11% | 28.89% |
@@ 3.7.2 Model comparison
The SVM model with linear kernel demonstrated the highest overall accuracy on the test data with an accuracy of 40.00 % and a mean accuracy of 39.02 % (SD = 5.50 %) during 10-fold cross-validation. In terms of class-specific performance on the test data, for a single class, the SVM model achieved balanced accuracies [11] of 70.68 % for anger, 87.65 % for attention, 70.68 % for happiness, and 72.53 % for calming. The model also performed relatively well for confusion with a balanced accuracy of 66.98 %. However, it struggled with classes such as disgust and fear, showing balanced accuracies of 60.49 % and 59.26 %, respectively. It performs worst in sadness and surprise with a balanced accuracy of 54.01 % and 54.63 %, respectively.
3.7.3 Confusion matrix
@@
The confusion matrix shown in Fig. 4 indicates that attention is the most recognisable among all emotions. However, several misclassifications are evident: anger is frequently misclassified as attention, comfort is often confused with calming, and happiness is mistakenly identified as attention. Additionally, sadness is commonly confused with both comfort and calming, while surprise is often misclassified as fear and attention.
@@
These misclassifications can be attributed to the overlapping characteristics of these emotions in the context of touch and its sound. For example, the physical sensation and auditory feedback associated with anger can resemble the intense, focused signals interpreted as attention. Additionally, when participants try to convey attention to the robot, the maximum pressure applied is greater than that used for fear, happiness, and confusion. This indicates that people might not be mindful of the applied force when attempting to grab the attention of a robot. Calming and comfort are easily confused due to their similar low arousal. Happiness, conveyed through touch and sound, might be mistaken for the engaged and focused nature of attention, and it is easily confused with surprise due to their similarly high arousal. Sadness is often confused with similarly low arousal states of comfort and calming. Surprise and fear all belong to high arousal emotions, making them difficult to distinguish from each other.
4 Conclusion
@@
One of our primary objectives was to determine if people consistently express the same emotions through touch. All the emotions showed significant consistency, while the ICC indicated varying degrees of consistency across different emotions. “Attention” achieved the highest consistency, suggesting that participants were more uniform in conveying this emotion. Conversely, “Surprise” had the lowest ICC, indicating considerable variability in how participants expressed this emotion through touch. The classification accuracy of the multi-modal also aligns with the ICC trend. The class “Attention” obtained the highest accuracy of 87.65 % while the class “Surprise” got the accuracy of 54.63%.
Regarding the distinguishability of emotions based on touch, consistency analyses indicated all the emotions are statistically significant consistency, although some emotions showed low ICC values. Furthermore, PERMANOVA results confirmed that most emotion pairs were significantly different. However, certain emotion pairs, like “Calming” and “Sadness” or “Confusion” and “Disgust”, did not exhibit significant differences. This suggests that these emotions share similar tactile and auditory characteristics, making them harder to distinguish.
Finally, we build a multi-modal to decode emotion; from the SVM model’s confusion matrix, we can conclude that emotions are easily misclassified into categories that share similar levels of arousal or valence.
The subjective feedback collected from participants revealed that surprise and confusion were the most challenging emotions to express, with 64.3% and 46.4% of participants, respectively, indicating these emotions as difficult to express based on touch. This feedback aligns with the lower consistency scores observed for these emotions. However, this study’s generalizability may be impacted by the limited sample size of 28 participants. Furthermore, the tactile sensor is currently limited to use only on the forearm.

### v2

@@ III-B Consistency analysis
In order to transform feature values to a similar scale, ensuring all features contribute equally to the model, we standardized the tactile and audio features. After standardization, the numeric features of the dataset have a mean of 0 and a standard deviation of 1. We then performed Principal Component Analysis (PCA) to reduce dimensionality and identify the key components capturing the most variance. The resulting principal component scores were integrated with the original emotion and participant identifiers. To evaluate the consistency of emotion conveyance, we calculated Intraclass Correlation Coefficients (ICC) for each emotion using the PCA scores and a two-way consistency model.
As shown in Fig. 5(a) and Table I, all emotions demonstrated statistically significant consistency across participants, although some emotions exhibited comparatively low ICC values. The highest ICC was observed for “Attention”, indicating the most consistent expression, while “Surprise” had the lowest ICC, suggesting the greatest variability in expression.
Fig. 5(b) and Table II demonstrate that all gestures exhibit significant consistency across participants. The “Hold” gesture achieved the highest ICC at 0.64, whereas the “Pat” gesture had the lowest ICC at 0.14. The “Tap” and “Tickle” gestures showed similar ICC values, at 0.35 and 0.31, respectively.
@@
| Emotions | Hold | Pat | Poke | Rub | Tap | Tickle |
| ICC | 0.64, | 0.14, | 0.16, | 0.19, | 0.35, | 0.31, |
III-C Variability analysis
III-C1 Variability analysis emotional touch expression
The features do not follow a normal distribution. Therefore, we used permutational multivariate analysis of variance (PERMANOVA), which reveals a significant effect of emotion on the scaled combined features (). Pairwise comparisons with Holm correction show that most emotion pairs differ significantly and have consistent significance after p-value adjustment (). However, there is no difference between “Calming” and “Sadness” (), “Comfort” and “Sadness” () based on touch and audio features, and there is no difference between “Confusion” and “Comfort” (), “Confusion” and “Sadness” (), “Disgust” and “Sadness” (), “Disgust” and “Surprise” (), “Happiness” and “Surprise” (). We hereby analyse the four valence and arousal quadrants plus one neutral emotion “Attention”, and the pairwise comparison with post-hoc analysis shows that all the pairwise comparison has significant differences between each other () except for Quadrant 2 and Quadrant 3 (), which is close to the significant level.
Intuitively, it is expected that “Surprise” is very similar to “Happiness”. One explanation could be that both of the emotions belong to the high-arousal zone. Similarly, there is no significant difference between “Confusion” and “Sadness”, or between “Disgust” and “Sadness”, as they belong to the low-arousal zone.
III-C2 Variability analysis for tactile gesture
Following the same procedure in affective touch variability analysis, refer to Subsection III-C1, the features do not follow a normal distribution. Therefore, we used PERMANOVA, which reveals a significant effect of gesture on the scaled combined features (). Pairwise comparisons with Holm correction show that all the gesture pairs differ significantly and have consistent significance after p-value adjustment ().
IV Decoding emotions using different modalities
IV-A Data split
@@
IV-B Model types
We used various machine learning techniques to decode the emotions conveyed by touch and sound.
IV-B1 Classical machine learning models
We selected Random Forest (RF) [38], Support Vector Machine (SVM) [39], and Decision Tree (DT)[40] for their robustness and proficiency in handling high-dimensional data. Specifically, RF excels at handling high-dimensional data and has built-in feature importance for classification [38, 41]. Decision trees are generally well-suited for small datasets and have been used for multimodal classification in [42, 43]. SVM has been used in touch gesture classification and audio emotion classification [44, 45]. The models were evaluated using 10-fold cross-validation to ensure robustness and accuracy. The results on the test dataset of these comparisons are presented in Table III. For all traditional models, we performed a grid search to optimize key hyperparameters, such as the number of estimators and maximum depth for RF, kernel type and regularization parameter for SVM, and depth for DT, using 10-fold cross-validation to ensure generalizability and prevent overfitting. All model architectures, training code, and hyperparameters are publicly available in our GitHub repository444https://github.com/qiaoqiao2323/Convey-emotion.
IV-B2 Deep learning models
In addition, deep learning is frequently used for emotion and touch gesture decoding. We explored five different models: a multi-temporal resolution convolutional neural network [46] (MTRCNN), a pretrained audio neural network (PANN) [47], and several models which have already been used before for touch gesture classification, namely a Convolution Neural Network - Gated Recurrent Unit (CNN_GRU), a CNN Transformer [48], and a CNN LSTM.
Each model architecture is described below.
MTRCNN is a multi-task convolutional network, using a multi-branch architecture with different convolution kernel sizes (3×3, 5×5, 7×7). Each branch includes a convolutional block followed by dilated convolutions and pooling layers to extract multi-scale features. The outputs of these branches are concatenated and passed through a fully connected embedding layer, followed by a classification layer for the final predictions.
CNN_LSTM integrates convolutional layers with an LSTM to capture both spatial and temporal features. It uses three convolutional layers followed by a two-layer LSTM to model temporal dependencies, and the final output is obtained through a fully connected layer.
CNN_Transformer combines convolutional feature extraction with Transformer-based temporal modelling. The input is passed through three convolutional layers, and then processed by a Transformer encoder to capture long-range dependencies, and the final output is obtained through a fully connected layer.
PANNs consists of six convolutional blocks, each progressively increasing feature dimensions. The network employs batch normalization, ReLU, dropout, and pooling within each block. After feature extraction, global frequency averaging and time aggregation are performed, followed by a fully connected layer for final class predictions.
The CNN_GRU model begins by applying three consecutive 3×3 convolutional layers. The number of output channels increases between layers, from 64 to 128, followed by 256. The GRU consists of two layers with a hidden size of 64 units, capturing the temporal dependencies in the input sequence. The GRU’s output is flattened and passed through a fully connected layer to make the class predictions.
IV-C Model training
Each model is trained on sound features only, on tactile features only, and on both feature sets combined, to examine whether multimodal emotion decoding outperforms single-modality decoding. We compared model fusion and feature fusion techniques and found that feature fusion generally yielded better performance. This could be due to the increased complexity introduced by model fusion for our relatively small dataset. Consequently, we opted for feature fusion for multimodal decoding. Results are averaged over 10 training runs.
IV-C1 Classical machine learning models
We use the Adam optimiser [49], with a default learning rate of 0.0001. Each model trains for 100 epochs.
IV-C2 Deep learning models
NN models were trained for 100 epochs using a batch size of 32. Performance evaluations were conducted at the end of each epoch. The optimisation employed a standard configuration without AdamW, using a fixed learning rate of 0.001 throughout training.
IV-D Decoding results & discussion
The results for gesture classification and emotion classification are given in Tables III and IV, respectively.
IV-D1 Emotion and touch gesture decoding
| Models test accuracy | MTRCNN | PANNs | CNN_GRU | CNN_Transformer | CNN_LSTM | RF | SVM | DT |
| Multimodal feature fusion | 81.52% | 81.24% | 82.54% | 76.21% | 90.74% | 81.85% | 80.56% | 73.15% |
| Sound model | 78.89% | 58.33% | 67.13% | 53.61% | 67.04% | 51.85% | 62.96% | 49.07% |
| Touch model | 79.53% | 80.97% | 81.39% | 74.59% | 79.63% | 79.54% | 77.78% | 72.22% |
| Models test accuracy | MTRCNN | PANNs | CNN_GRU | CNN_Transformer | CNN_LSTM | RF | SVM | DT |
| Multimodal feature fusion | 31.11% | 29.61% | 40.00% | 28.89% | 31.85% | 38.44% | 40.00% | 31.11% |
| Sound model | 24.72% | 23.78% | 28.33% | 17.44% | 26.17% | 23.00% | 22.55% | 23.89% |
| Touch model | 23.33% | 28.95% | 22.00% | 20.74% | 20.31% | 36.83% | 36.11% | 30.28% |
@@ IV-D1 Emotion and touch gesture decoding
For the test data, both the SVM model with a linear kernel and the CNN_GRU model showed the highest overall accuracy, each achieving 40.00 % in multimodal emotion decoding, which is significantly higher than the chance level of 10 % (p<0.05). Specifically, the SVM model scored 40% accuracy when combining sound and touch modalities, 22.55 % with just sound, and 36.11 % with just touch. The CNN_GRU model also reached 40% accuracy in the dual modality setup but achieved lower scores of 28.33 % and 22.00 % with sound and touch modalities, respectively. In terms of class-specific performance on the test data, for a single class, the SVM model achieved balanced accuracies [50] of 70.68 % for anger, 87.65 % for “Attention”, 70.68 % for “Happiness”, and 72.53 % for “Calming”. The model also performed relatively well for “Confusion” with a balanced accuracy of 66.98 %. However, it struggled with classes such as “Disgust” and “Fear”, showing balanced accuracies of 60.49 % and 59.26 %, respectively. It performs worst in “Sadness” and “Surprise” with a balanced accuracy of 54.01 % and 54.63 %, respectively.
Previous research [51, 4] explored emotion decoding among couples and strangers, and the results we obtained in our research are similar to the results they presented for strangers (37.5%). If we set the 37.5% as a baseline, we used a one-sample t-test to determine whether the participant’s decoding accuracy is significantly higher than chance level; the results showed that there is no significant difference between human-to-human emotion decoding among strangers in previous research and our human-robot decoding ().
Based on the comparison results in Table III and Table IV, we observed that the tactile-sound multimodal emotion recognition model achieved an accuracy of 40 %, demonstrating an improvement of approximately 3.89 % over the touch modality and 11.67 % over the sound modality alone. For touch gesture decoding, the tactile-sound model also showed a notable improvement, with an accuracy increase of 11.85 % compared to the sound modality and 9.35 % compared to the touch modality. These results highlight the effectiveness of integrating both tactile and auditory information, suggesting that combining the two modalities enhances the system’s ability to decode both emotions and gestures more accurately than using either modality alone.
We observed that emotion decoding based on the sound modality performed worse than the touch modality. This discrepancy could be due to the fact that participants were required to express emotions primarily through touch expressions, leading them to focus on tactile features such as amplitude, intensity, and contact area. The sound generated by touch may not fully capture the richness of these tactile expressions.
For instance, a firm grasp typically involves high force and a large contact area, which is well-detected by the tactile sensor but only produces minimal auditory feedback. Similarly, a high-speed shaking motion, which many participants used to express fear, generates significant tactile feedback but may result in weak or inconsistent sound recordings, failing to capture the full dynamics of the touch behaviour. This limitation in the sound modality could explain why participants found it more challenging to decode emotions based on auditory cues alone.
IV-D2 Affective touch confusion matrix
@@
The confusion matrix shown in Fig. 6 indicates that “Attention” is the most recognisable among all emotions. However, several misclassifications are evident: “Anger” is frequently misclassified as “Attention”, “Comfort” is often confused with “Calming”, and “Happiness” is mistakenly identified as “Attention”. Additionally, “Sadness” is commonly confused with both “Comfort” and “Calming”, while “Surprise” is often misclassified as “Fear” and “Attention”.
@@
These misclassifications can be attributed to the overlapping characteristics of these emotions in the context of touch and its sound. For example, the physical sensation and auditory feedback associated with “anger” can resemble the intense, focused signals interpreted as “Attention”. Additionally, when participants try to convey “Attention” to the robot, the maximum pressure applied is greater than that used for “Fear”, “Happiness”, and “Confusion”. This indicates that people might not be mindful of the applied force when attempting to grab the attention of a robot. “Calming” and “Comfort” are easily confused due to their similar low arousal. “Happiness”, conveyed through touch and sound, might be mistaken for the engaged and focused nature of “Attention”, and it is easily confused with “Surprise” due to their similarly high arousal. “Sadness” is often confused with similarly low arousal states of “Comfort” and “Calming”. “Surprise” and fear all belong to high arousal emotions, making them difficult to distinguish from each other.
IV-D3 Tactile gestures confusion matrix
The overall performance of the tactile gestures surpasses that of the affective touch, which achieved 90.74%. As shown in Fig. 7, the “Pat” gesture is once misclassified as “Tap”, while “Poke” is misclassified twice to “Tap” and “Pat”, respectively. In addition, “Rub” has sometimes been misclassified as “Tickle” and “Hold” and “Hold” sometimes misclassified to “Pat”, “Poke” and “Tap” are quite similar, especially when participants perform a quick poke to the robot. Additionally, “Tickle” and “Tap” are confused, which might be due to the similar motion and contact pattern involved in these gestures. “Rub”, “Hold”, and “Pat” are confused with each other, which could be due to the similar touching area of both gestures. In conclusion, the misclassification likely arises from the similar feature in execution and the overlapping sensory signal these gestures provide.
V Discussion
@@
In this research, we built tactile sensing capability for the Pepper robot by designing a piezoresistive pressure sensor. One of our primary objectives was to determine whether people consistently express the same emotions and touch gestures through touch. All the emotions showed significant consistency, while the ICC indicated varying degrees of consistency across different emotions. “Attention” achieved the highest consistency, suggesting that participants were more uniform in conveying this emotion. Conversely, “Surprise” had the lowest ICC, indicating considerable variability in how participants expressed this emotion through touch. The classification accuracy of the multimodal also aligns with the ICC trend. The class “Attention” obtained the highest accuracy of 87.65 % while the class “Surprise” got the accuracy of 54.63%. Similarly, all touch gestures demonstrated significant consistency, with ICC values for touch gestures generally higher than those for emotions, except for “Attention”. This suggests that people have more consensus on touch gestures than on emotions. One possible explanation is that, although emotions correlate with touch gestures, they can be expressed through various touch gestures. Additionally, people’s perception of emotions can be conveyed through different touch patterns and features. In contrast, social touch gestures often have a common understanding. We also found that “Attention” has a higher correlation with certain gestures, such as “Tap” or “Poke”, leading to higher ICC values among all emotions. This consistency in specific gestures for “Attention” could explain the higher ICC values and classification accuracy.
The phenomenon of varying consistency in touch gestures and emotions can be attributed to several factors. Touch gestures are often straightforward and universally understood, such as a tap or poke to convey attention, leading to higher consistency across individuals. In contrast, emotions are more complex and can be expressed through a variety of touch gestures. For example, love can be shown through a hug, a gentle touch, or a pat, resulting in lower consistency as different participants have different strategies to express different emotions. Previous research asked participants [17] to convey four emotions using specific gestures like tapping, rubbing, and pressing. In our research, participants were not given a specific list of gestures, leading to more variability in touch-based expression. Additionally, cultural norms and personal habits significantly influence how people express emotions through touch, with different cultures having varying interpretations and norms. Even though all participants in our study were from the same culture, personal experiences and comfort levels with touch also contributed to this variability. The context of interaction further affects how touch is interpreted, as a touch meant to convey attention in a formal setting might differ from one in a casual or intimate setting. The inherent complexity and multifaceted nature of emotions mean that a single touch gesture might not fully capture an emotion, leading to a broader range of expressions. The higher consistency in touch gestures compared to emotions is thus due to their universal nature, while the complexity and variability of emotions, influenced by cultural, personal, and contextual factors, lead to greater differences in expression.
Our findings indicate that touch gestures are generally more distinguishable than emotions when based on tactile interaction. The PERMANOVA results confirmed that all gestures were significantly different from each other, suggesting strong consensus among participants in their execution. Gestures like “Tap” and “Tickle” were among the easiest to distinguish due to their unique characteristics— “Poke” being a short, localized, high-intensity action and “Rub” involving continuous motion over a larger area. “Tap” was also relatively easy to decode due to its brief and repetitive nature. However, “Pat” and “Tap” were sometimes confused, likely due to similarities in their intensity and contact area. “Pat”, in particular, shares overlapping features with both “Tap” and “Tickle”, making it more prone to misclassification. In contrast, emotions were more challenging to differentiate, particularly when they shared similar tactile and auditory characteristics. Although all emotions showed statistically significant consistency, certain pairs, such as “Calming” and “Sadness” or “Confusion” and “Disgust”, did not exhibit significant differences. This suggests that these emotions share common touch patterns, making them harder to distinguish. “Happiness” and “Anger” were among the easiest to decode, as they were often associated with distinct tactile expressions, “Happiness” frequently involved rhythmic, gentle movements, while “Anger” was characterized by stronger pressure and abrupt interactions. “Fear” and “Disgust” were moderately distinguishable but sometimes confused due to their overlapping aversive touch patterns.
The confusion matrix from the SVM model further revealed that emotions were often misclassified into categories with similar arousal or valence levels. For instance, emotions in the low-arousal range, such as “Calming” and “Sadness”, were frequently mixed up, while high-arousal emotions like “Surprise” and “Fear” also showed overlap. Subjective feedback supported these findings, with “Surprise” and “Confusion” being the most difficult emotions to express through touch, reported by 64.3% and 46.4% of participants, respectively. Overall, gesture classification achieved higher accuracy than emotion classification, reinforcing the idea that social touch gestures are more consistently performed and interpreted. While some emotions can be effectively conveyed through touch, those with similar arousal or valence levels present greater challenges for clear differentiation. These findings highlight the complexity of affective touch and suggest that while gestures provide structured cues for communication, emotional expression through touch requires additional contextual or multimodal cues for enhanced recognition.
Our study confirms that multimodal models integrating touch and sound features outperform single-modal approaches in decoding emotions and gestures, achieving higher classification accuracy and robustness. The tactile-sound dual model improved emotion recognition by 3.89% over the touch-only model and 11.67% over the sound-only model, while also enhancing gesture decoding by 11.85% and 9.35% over sound and touch modalities, respectively. These improvements stem from the complementary nature of touch and sound cues, where tactile features capture intensity, amplitude, and contact area, while audio signals add temporal dynamics, compensating for the weaknesses of each modality. Single-modal sound decoding, for instance, struggled with capturing high-force gestures like firm grasps or high-speed shaking, leading to information loss. By fusing modalities, multimodal models provide richer, more diverse feature representations, making them more effective in recognizing subtle emotional and gestural variations. This has significant implications for human-robot interaction, particularly in social, assistive, and affective computing applications, where natural and adaptive emotional communication is essential. Future research should explore expanding sensor coverage, incorporating additional modalities (e.g., physiological signals), and improving context-awareness to further enhance robotic affective touch systems.
V-A Limitations
However, the generalizability of this study may be limited by the relatively small sample size of 28 participants, all from a single cultural background. Additionally, the current tactile sensor is restricted to the forearm, which may not capture the full range of touch expressions used in natural interactions. Although emotions can be conveyed through touch alone, our findings show that it remains difficult to distinguish between emotions with similar arousal and valence characteristics based solely on tactile signals. In real-world scenarios, human emotional communication is inherently multimodal, involving visual cues (e.g., facial expressions, body posture) and verbal elements (e.g., tone, content of speech). The absence of these complementary modalities in our study may have limited participants’ ability to express emotions naturally and constrained the system’s decoding accuracy. Furthermore, the interpretation of touch gestures is highly context-dependent—not only influenced by the physical design of the robot but also by its perceived social role. A robot’s appearance, including its level of human-likeness and perceived warmth, can affect users’ comfort and willingness to engage in tactile interaction. For example, a more human-like and friendly-looking robot may encourage more natural and expressive touch gestures than a mechanical or utilitarian one. Similarly, the situational context—such as whether the robot is perceived as a caregiver, companion, or assistant—can shape both the types of touch used and how those touches are interpreted. These factors introduce ambiguity into the decoding process and underscore the importance of designing tactile communication systems that are sensitive to both social context and the robot’s embodied characteristics.